[figure: 9.7 cm]

SPONTANEOUS AND MAGNETIC FIELD INDUCED PHASE

TRANSITIONS IN FERROELECTRIC Y Mn_{2}O_{5}

Yu.F. Popov^{a}, A.M. Kadomtseva^{a}, G.P.Vorob’ev^{a}, K.I. Kamilov^{a}, M.M. Lukina^{a}, Mahov P.N.^{a} and M.M. Tegranchi^{b} aMoscow State University, 119992 Moscow, Russian Federation bDepartment of Physics, Shahid Beheshti University, Evin, 19839 Tehran

Magnetic, magnetoelectric and magnetoelastic properties of Y Mn_{2}O_{5} single crystals were investigated in static (up to 12 kOe) and pulse (up to 250 kOe) magnetic fields. Exsitance in Y Mn_{2}O_{5} below T_{N} = 42K of the spontaneous weak ferromagnetic moment along a- axis of rhombic crystal that determined by magnetoelectric interaction was experimentally found and theoretically approved. At T_{C} = 20K in Y Mn_{2}O_{5} ferroelectric transition has been observed with emerge of spontaneous electric polarization along baxis of crystal[1] . It was shown that at T=25 K magnetic field along a- axis of crystal at H_{C} = 100kOe call sharp jump of electrical polarization along b- axis that can be explained by increasing of temperature of ferroelectric transitin. It was found that at strong enough magnetic field H——b in P_{b}(H_{b}) dependence jumps of polarization have been emerged. It explained by electric polazition sign changing at reorientation of magnetic moment. Reorientation of spins accompanied by magnetostriction deformities. This work is support by RFBR (00-15-96695 and 01-02-164486)

1. A. Ikeda, K. Kohn, Ferroelectrics, 169 (1995) 75.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : K.I. Kamilov

Address for correspondence : Department of Physics Moscow State University Leninskie Gori, 119992 Moscow, Russian Federation

Email address : kamilov@plms.phys.msu.su